# A Thermodynamically Consistent Free Boundary Model for Two-Phase Flows in an Evolving Domain with Bulk-Surface Interaction

### v2

@@ 3.3  Completion of the model derivation by the Lagrange multiplier approach
$\quad-\mu\big(\partial_{t}\phi+\textnormal{div}\,(\phi{\mathbf{u}})+\textnormal{div}\,\,\mathbf{J}_{\phi}\big)-p\,\textnormal{div}\,{\mathbf{u}}$ in $Q_{T}$, (3.50)
@@
$\quad-\theta\big(\partial_{t}^{\circ}\psi+\textnormal{div}_{\Gamma}\,\mathbf{K}_{\psi}-\mathbf{J}_{\psi}\cdot{\mathbf{n}}\big)-q\,\textnormal{div}_{\Gamma}{\mathbf{u}}$ on $S_{T}$. (3.51)
In the following, we will simply write $\rho$, $f$ and $g$ instead of $\rho(\phi)$, $f(\phi,\nabla\phi)$ and $g(\phi,\psi,\nabla_{\Gamma}\psi)$ to provide a cleaner presentation. By the definition of the energy densities $e_{\Omega}$ and $e_{\Gamma}$ (see (3.39) and (3.41)), we reformulate (3.3) and (3.3) as
@@
3.3.1 Computations in the bulk
In the bulk, we proceed exactly as in [5] and [50] to reformulate inequality (3.3) as
@@ 3.3.1 Computations in the bulk
Due to these choices, the first two lines of the right-hand side of (3.56) vanish. We further assume the mass flux $\mathbf{J}_{\phi}$ to follow Fick’s law. This means that
@@
where the nonnegative function $\nu=\nu(\phi)$ represents the viscosity of the fluids, and $\mathbf{D}{\mathbf{u}}$ denotes the symmetric gradient of $\mathbf{u}$}. This choice along with (3.57), (3.58) and the identity $p\mathbf{I}:\nabla{\mathbf{u}}=p\,\textnormal{div}\,\,{\mathbf{u}}=0$ in $Q_{T}$ ensures that (3.56) is fulfilled as it reduces to
@@
3.3.2 Computations on the boundary
We now consider the local energy dissipation law (3.3) on the surface. Recalling the formulas for $f$ (see (3.40)) and $\mathbf{J}_{e}$ (see (3.58)), we deduce from (3.3) that
@@ 3.3.2 Computations on the boundary
on $S_{T}$. Now, the most technical step is to reformulate the term $\partial_{\nabla_{\Gamma}\psi}g\cdot\partial_{t}^{\circ}\nabla_{\Gamma}\psi$. According to [17, Lemma 38], we have
@@
$\delta(\nabla_{\Gamma}\psi\otimes\nabla_{\Gamma}\psi){\mathbf{u}}\cdot{\mathbf{n}}=0$ (3.72)
@@
on $S_{T}$. We now assume that the flux terms $\mathbf{J}_{\phi}$ and $\mathbf{J}_{\psi}$ are directly proportional. This means that there exists $\beta\in\mathbb{R}$ such that
$\mathbf{J}_{\psi}\cdot{\mathbf{n}}=\beta\mathbf{J}_{\phi}\cdot{\mathbf{n}}=-\beta\,m_{\Omega}(\phi)\nabla\mu\quad\quad\text{on }S_{T}.$ (3.78)
Here, the second equality follows from (3.59). We further make the constitutive assumptions
@@
where $\gamma_{\tau}$ and $\gamma_{\mathbf{n}}$ are nonnegative functions, which may depend on the phase fields. Equation (3.79) can be interpreted as a generalized Navier slip boundary condition with $\gamma_{\tau}$ being the slip parameter. We point out that an analogous boundary condition appeared in [50]. By means of (3.78)–(3.80), inequality (3.77) further reduces to
@@
$\theta=-\delta\Delta_{\Gamma}\psi+\frac{1}{\delta}G^{\prime}(\psi)+\alpha\varepsilon\nabla\phi\cdot{\mathbf{n}}\quad\text{on }S_{T}.$ (3.84)
@@
In summary, we have thus derived system (1.1) along with formulas (1.2)-(1.4) with the following correspondences:

### v3

@@ 3.2.2 The Lagrange Multiplier Approach
$\quad-\mu\underbrace{\big(\partial_{t}\phi+\textnormal{div}\,(\phi{\mathbf{u}})+\textnormal{div}\,\,\mathbf{J}_{\phi}\big)}_{=0}-\,p\,\underbrace{\textnormal{div}\,{\mathbf{u}}\phantom{\big|\!}}_{=0}$ in $Q_{T}$, (3.51)
@@
$\quad-\theta\underbrace{\big(\partial_{t}^{\circ}\psi+\textnormal{div}_{\Gamma}\,\mathbf{K}_{\psi}-\mathbf{J}_{\psi}\cdot{\mathbf{n}}\big)}_{=0}-\,q\,\underbrace{\textnormal{div}_{\Gamma}{\mathbf{u}}\phantom{\big|\!}}_{=0}$ on $S_{T}$. (3.52)
In the following, we will simply write $\rho$, $f$ and $g$ instead of $\rho(\phi)$, $f(\phi,\nabla\phi)$ and $g(\phi,\psi,\nabla_{\Gamma}\psi)$ to provide a cleaner presentation. By the definition of the energy densities $e_{\Omega}$ and $e_{\Gamma}$ (see (3.40) and (3.42)), we reformulate (3.2.2) and (3.2.2) as
@@
Computations in the bulk.
In the bulk, we proceed exactly as in [Abels2012] and [Giorgini2023] to reformulate inequality (3.2.2) as
@@
Note that in (3.58), we could potentially add a term $\eta(\phi)\,\partial_{t}^{\bullet}\phi$ (with $\eta(\phi)>0$) on the right-hand side of (3.58) to account for dissipative friction. This would cause an additional dissipative term $\eta(\phi)\left|\partial_{t}^{\bullet}\phi\right|\leq 0$ on the right-hand side of (3.57), which is consistent with the local energy dissipation law. In the final system of equations, this term would then lead to a viscous relaxation of the Cahn–Hilliard equation. However, in our model, we assume dissipative friction to be negligible and therefore, we do not include such a term. Due to the choices (3.58) and (3.59), the first two lines of the right-hand side of (3.57) vanish. We further assume the mass flux $\mathbf{J}_{\phi}$ to follow Fick’s law. This means that
@@
where the nonnegative function $\nu=\nu(\phi)$ represents the viscosity of fluids, and $\mathbf{D}{\mathbf{u}}$ denotes the symmetric gradient of $\mathbf{u}$}. This choice along with (3.58), (3.59) and the identity $p\mathbf{I}:\nabla{\mathbf{u}}=p\,\textnormal{div}\,\,{\mathbf{u}}=0$ in $Q_{T}$ ensures that (3.57) is fulfilled as it reduces to
@@
Computations on the boundary.
We now consider the local energy dissipation law (3.2.2) on the surface. Recalling the formulae for $f$ (see (3.41)) and $\mathbf{J}_{e}$ (see (3.59)), we deduce from (3.2.2) that
@@
on $S_{T}$. Now, the most technical step is to reformulate the term $\partial_{\nabla_{\Gamma}\psi}g\cdot\partial_{t}^{\circ}\nabla_{\Gamma}\psi$. According to [BGN-book, Lemma 38], we have
@@
${\delta(\nabla_{\Gamma}\psi\otimes\nabla_{\Gamma}\psi){\mathbf{u}}\cdot{\mathbf{n}}=\delta(\nabla_{\Gamma}\psi\cdot{\mathbf{u}})(\nabla_{\Gamma}\psi\cdot{\mathbf{n}})=0}$ (3.73)
@@
on $S_{T}$. We now assume that the flux terms $\mathbf{J}_{\phi}$ and $\mathbf{J}_{\psi}$ are directly proportional. This means that there exists a coefficient $\beta$ such that
$\mathbf{J}_{\psi}\cdot{\mathbf{n}}=\beta\mathbf{J}_{\phi}\cdot{\mathbf{n}}=-\beta\,m_{\Omega}(\phi){\nabla\mu\cdot{\mathbf{n}}}\quad\quad\text{on }S_{T}.$ (3.79)
Here, the second equality follows from (3.60). The idea behind this assumption is that $\beta$ represents the fraction of bulk material that is effectively transformed into surface material when reaching the boundary. In this paper, we assume for simplicity that $\beta$ is a constant. However, in general, the coefficient could also be a sufficiently regular function $\beta:Q_{T}\to\mathbb{R}$ (see, e.g., [Goldstein2011]).
We further make the constitutive assumptions
@@
where $\gamma_{\tau}$ and $\gamma_{\mathbf{n}}$ are nonnegative functions, which may depend on the phase fields. Equation (3.80) can be interpreted as a generalized Navier slip boundary condition with $\gamma_{\tau}$ being the slip parameter. We point out that an analogous boundary condition appeared in [Giorgini2023]. The boundary condition (3.81) relates the normal component of $\mathbf{u}$} to the normal traction $\mathbf{T}{\mathbf{n}}$ and the geometry of the evolving boundary through its mean curvature $H$. The coefficient $\gamma_{\mathbf{n}}>0$ acts as a weighting parameter for the energy dissipation due to boundary deformation (cf. (1.10)). By means of (3.79)–(3.81), inequality (3.78) further reduces to
@@
$\theta=-\delta\Delta_{\Gamma}\psi+\frac{1}{\delta}G^{\prime}(\psi)+\alpha\varepsilon\partial_{\mathbf{n}}\phi\quad\text{on }S_{T}.$ (3.85)
@@
In summary, we have thus derived system (1.1) along with formulae (1.2) and (1.4) with the following correspondences: